Calculate the inductive reactance in ohms of a single phase, 8 kV, 50 hz, 50 km long transmission line consisting of two aluminium conductors with a 3 cm diameter by a spacing of 2 meters. Then, calculate the shunt admittance of the above line in siemens per km.
Answer #1

Answer:

Given that:

Calculate the inductive reactance in ohms of a single phase , 8 kV, 50 hZ, 50 km long transmission

line consisting of two aluminium conductors with a 3 cm diameter by a spacing of 2 meters .

a) inductive reactance :

two series inductance per meter of two transmission line is

= 2.057*10-6 H/M

inductance of the transmission line will be

l =(2.057*10-6 ) (50* 1000)

= 0.1029H

AC resistance of line will be

RAC = (2) (1.05)

= 2.1

Total series impedance of this line would be

Z = 2.1 + j 30.5

impedance / kM will be

Z = (2.1 + j 32.3) / 50

Z = (0.042 + 30.646 ) /kM

Shunt admittance of the above line in siemens per km.

The shunt capacitance /meter of this transmission line is

Capacitance / kM = 5.69* 10-9 F/M

Shunt admittance   Y 8h = j2fc = (j1.79* 10-6) g / kM

total shunt admittance Y 8h = (1.79*10-6 ) (50*103) = 8.95*10-5